# Manipulating Spectral Windings and Skin Modes through Nonconservative Couplings

### v1

@@ Abstract
Connection with nonreciprocity.— Extending the analysis of the connection between the nonreciprocity performance of the chain and the NHSE shown in different phase regions, we consider two different nonreciprocal transmission cases by choosing the leftmost mode $a_{1}$ ($b_{1}$) and rightmost mode $a_{N}$ ($b_{N}$) to connect the input and output field, respectively. Therefore, the nonreciprocity ratios of the two cases can be defined as $I_{a}=T_{a\leftarrow}/T_{a\rightarrow}=|S_{1,2N-1}/S_{2N-1,1}|^{2}$, $I_{b}=T_{b\leftarrow}/T_{b\rightarrow}=|S_{2,2N}/S_{2N,2}|^{2}$, which can be calculated by solving the system scattering matrix $S$ numerically (see the Supplemental Material for more details [46]). Once the off-diagonal elements of $S$ satisfy $|S_{1,2N-1}|\neq|S_{2N-1,1}|$ ($|S_{2,2N}|\neq|S_{2N,2}|$), the nonreciprocal transmission in the chain can be achieved. When the transmission efficiency for the backward (forward) direction is larger than that for the forward (backward) direction for the chain, i.e., $I_{a/b}>1$ ($I_{a/b}<1$), the nonreciprocity direction of the chain is on the backward (forward).
Figure 4: Diagrams of nonreciprocity ratios as functions of the nonconservative coupling phases $\phi_{a}$ and $\phi_{b}$, with the input/output field connected to mode $a_{1}/a_{N}$(a) and $b_{1}/b_{N}$(b). The red indicates the nonreciprocity ratio $I_{a/b}>1$, and the blue region indicates $I_{a/b}<1$. The size of the chain is $N=4$ and the other parameters are the same as those presented in Fig. 2.
@@
The nonreciprocity ratios of the two different settings that the input/output field connecting with the mode $a_{1}$/$a_{N}$ and $b_{1}$/$b_{N}$ are illustrated in Fig. 4 (a) and (b), respectively. The red (blue) region represents the nonreciprocity ratio $I_{a/b}>1$ ($I_{a/b}<1$), indicating the backward (forward) nonreciprocity direction for the chain. Comparing the spectral phase diagram shown in Fig. 2 (a) and (b), we can find the nonreciprocal transmission shows similar behaviors with the eigenmode localization of NHSE. When tuning the phases as $\text{sgn}(\phi_{a})=-\text{sgn}(\phi_{b})<0$ and $|\phi_{a}|$ and $|\phi_{b}|$ are around $\pi/2$, both $I_{a}$ and $I_{b}$ are greater than 1 for left eigenmode localization and less than 1 for right eigenmode localization of unipolar NHSE. However, the nonreciprocity directions of the two input/output settings will become opposite when tuning the nonconservative coupling phases to the region of closed loops with opposite windings. We can find that the nonreciprocity ratio $I_{a}>1$ while $I_{b}<1$ when $\text{sgn}(\phi_{a})=\text{sgn}(\phi_{b})>0$, and $I_{a}<1$ while $I_{b}>1$ when $\text{sgn}(\phi_{a})=\text{sgn}(\phi_{b})<0$ [Fig. 4 (a) and (b)]. Therefore, the opposite nonreciprocity directions of the two settings can connect with the opposite localization of the two groups of eigenmodes, indicating a close link between the bipolar NHSE and nonreciprocity. In addition, when approaching the band twisting region, the change of the nonreciprocity direction is determined by the values of the nonconservative coupling phases and the chain size $N$ together (see the Supplemental Material for more details [46]).
@@
Conclusion.— In conclusion, we elucidated the manipulation of the NHSE through nonconservative couplings that extend beyond nearest-neighbor couplings, demonstrating multi-type spectral phases and their impact on mode localizations. By tuning the phases of the nonconservative couplings, we achieved diverse spectral structures associated with both unipolar and bipolar NHSE localizations. Moreover, we have evidenced that multi-type spectral phases as well as NHSE are closely linked to the nonreciprocal transmission of the chain. Our approach is also experimentally feasible as nonconservative couplings have been realized in various systems, including optical systems [48, 49], room-temperature atomic ensembles [50] and magnonics system [51, 52]. Our findings not only provide new insights into the control of NHSE but also pave the way for advanced applications in controlling nonreciprocal energy flow.

### v2

@@ IV Connection with nonreciprocity
Extending the analysis of the connection between the nonreciprocity performance of the chain and the NHSE shown in different phase regions, we consider two different nonreciprocal transmission cases by choosing the leftmost mode $a_{1}$ ($b_{1}$) and rightmost mode $a_{N}$ ($b_{N}$) to connect the input and output field, respectively. Therefore, the nonreciprocity ratios of the two cases can be defined as $I_{a}=T_{a\leftarrow}/T_{a\rightarrow}=|S_{1,2N-1}/S_{2N-1,1}|^{2}$, $I_{b}=T_{b\leftarrow}/T_{b\rightarrow}=|S_{2,2N}/S_{2N,2}|^{2}$, which can be calculated by solving the system scattering matrix $S$ numerically (see the Appendix C for more details). Once the off-diagonal elements of $S$ satisfy $|S_{1,2N-1}|\neq|S_{2N-1,1}|$ ($|S_{2,2N}|\neq|S_{2N,2}|$), the nonreciprocal transmission in the chain can be achieved. When the transmission efficiency for the backward (forward) direction is larger than that for the forward (backward) direction for the chain, i.e., $I_{a/b}>1$ ($I_{a/b}<1$), the nonreciprocity direction of the chain is backward (forward).
@@
The nonreciprocity ratios of the two different settings for the input and output field connecting with the mode $a_{1}$/$a_{N}$ and $b_{1}$/$b_{N}$ are illustrated in Fig. 6 (a)(c) and (b)(d), respectively. The red (blue) region represents the nonreciprocity ratio $I_{a/b}>1$ ($I_{a/b}<1$), indicating the backward (forward) nonreciprocity direction for the chain. Comparing the spectral phase diagram shown in Fig. 2 (a) and (b), we can find the nonreciprocal transmission shows similar behaviors with the eigenmode localization of NHSE. When tuning the phases as $\text{sgn}(\phi_{a})=-\text{sgn}(\phi_{b})<0$ and $|\phi_{a}|$ and $|\phi_{b}|$ are around $\pi/2$, both $I_{a}$ and $I_{b}$ are greater than 1 for left eigenmode localization and less than 1 for right eigenmode localization of unipolar NHSE. However, the nonreciprocity directions of the two input and output settings will become opposite when tuning the nonconservative coupling phases to the region of closed loops with opposite windings. We can find that the nonreciprocity ratio $I_{a}>1$ while $I_{b}<1$ when $\text{sgn}(\phi_{a})=\text{sgn}(\phi_{b})>0$, and $I_{a}<1$ while $I_{b}>1$ when $\text{sgn}(\phi_{a})=\text{sgn}(\phi_{b})<0$ [Fig. 6 (a)(c) and (b)(d)]. Therefore, the opposite nonreciprocity directions of the two settings can connect with the opposite localization of the two groups of eigenmodes, indicating a close link between the bipolar NHSE and nonreciprocity.
Moreover, as we approach the band twisting region, the change of the nonreciprocity direction is determined by the values of the nonconservative coupling phases and the chain size $N$ together. As illustrated in Fig. 6, comparing the Fig. 6(a)(b) with (c)(d), which represent different system sizes, we observe that the change in nonreciprocity direction is influenced by both the nonconservative coupling phases and the chain size $N$. Due to the rapid decrease in transmission efficiency with increasing $N$, we limit our analysis of nonreciprocity to chains where $N\leq 5$.
Figure 6: Diagrams of nonreciprocity ratios as functions of the nonconservative coupling phases $\phi_{a}$ and $\phi_{b}$, with the input and output field connected to mode $a_{1}/a_{N}$(a)(c) and $b_{1}/b_{N}$(b)(d) with different sizes of the chain. The red indicates the nonreciprocity ratio $I_{a/b}>1$, and the blue region indicates $I_{a/b}<1$. The size of the chain is $N=4$(a)(b) and $5$(c)(d) and the other parameters are $w/\delta=1/2,\,u/\delta=i/2,\,v/\delta=1/2$.
V Conclusion
@@
In conclusion, we elucidated the manipulation of the NHSE through nonconservative couplings that extend beyond nearest-neighbor couplings, demonstrating multi-type spectral phases and their impact on mode localizations. By tuning the phases of the nonconservative couplings, we achieved diverse spectral structures associated with both unipolar and bipolar NHSE localizations. Moreover, we have evidenced that multi-type spectral phases as well as NHSE are closely linked to the nonreciprocal transmission of the chain. Our approach is also experimentally feasible as nonconservative couplings have been realized in various systems, including optical systems [50, 51], room-temperature atomic ensembles [52] and magnonics system [53, 54]. Our findings not only provide new insights into the control of NHSE but also pave the way for advanced applications in controlling nonreciprocal energy flow.
Acknowledgements.